OpenStudy (anonymous):

the equation of the common tangent to the curves y^2 = 8x and xy = -1 is?

OpenStudy (anonymous):

@Astrophysics can you help me out?

OpenStudy (anonymous):

so I take a general tangent to one of the curves, then I apply the condition that it is tangent to the other.

OpenStudy (anonymous):

as a matter of fact I know that the general tangent of y^2 = 8x is y = mx + 2/m

OpenStudy (anonymous):

xy = -1 now I've got y' = 1/x^2 what next?

OpenStudy (astrophysics):

Ok good so we have \[y=mx+\frac{ 2 }{ m }\] for \[y^2 = 8x\] then your second equation \[xy=-1\] should have the tangent. \[x \left( mx+\frac{ 2 }{ m } \right) = -1\] find the m

OpenStudy (astrophysics):

You can use differentiation if you like but I don't think you need to

OpenStudy (anonymous):

you wrote x(mx+2/m) = -1 as the tangent for the second curve but that's not even a line unless I'm missing something like if m is dependent on x.

OpenStudy (astrophysics):

No I mean the tangent also touches the curve xy=-1

OpenStudy (anonymous):

oh of course yeah, that makes sense so how would we go about that?

OpenStudy (astrophysics):

\[x \left( mx+\frac{ 2 }{ m } \right) = -1 \] solving the quadratic will give you the value of your slope here which will give you the tangent for BOTH of these equations once you plug the values in \[y=mx+\frac{ 2 }{ m } \]

OpenStudy (anonymous):

will that really work?

Parth (parthkohli):

Use discriminants! That quadratic equation will only have one solution. So equate the discriminant to zero.